a vehicle travels from A to B in 8 sec. at a constant speed of 25 mph. A second vehicle travels the distance in 8 sec. What is the distance from A to B and how fast is the second vehicle traveling ?
1
Expert's answer
2012-05-25T08:53:32-0400
v=25, t=8/3600=1/450 => s=vt=25/450=1/9=0.111111111.....
So, speed of
second vehicle v'=s/(8 sec)=25 mph